# Performance Analysis and Power Allocation for Massive MIMO ISAC Systems

### v2

@@ III-B Sensing Performance
See Appendix D. $\Box$
We have shown in Remarks 2 and 3 that with the simplified equal power allocation in (23), the interference is mitigated and a very low CRLB can be achieved by a very large $N_{\mathtt{t}}$. However, this does not guarantee a good communications–sensing performance tradeoff. In particular, Remark 1 implies that the massive MIMO ISAC system requires a proper power allocation between communications and sensing; otherwise, there will be significant performance loss due to sensing. In the next section, we present the power allocation problem and its solution.
@@ V-C Communications and Sensing Performance
In this section, we demonstrate the communications and sensing performance of MRT and ZF precoding with the proposed power allocation. We first verify the theoretical results for the CRLB in Theorems 1 and 2. In Fig. 3, we compare the CRLB for $\theta$ and $\phi$ obtained by Monte-Carlo simulations based on (37), (38), and the analytical results in Theorem 2. We also show the MSE of the MLE in (27) for $\theta$ and $\phi$, which performs an exhaustive search over a fine grid. The sensing SNR is defined as $\frac{P_{\mathtt{t}}L\left|\alpha\right|^{2}}{\sigma_{\mathtt{s}}^{2}}$ [22]. It is observed for both MRT and ZF that the CRLB is a tight lower bound for the MSE of MLE. Therefore, the power allocation constraint on the CRLB threshold can guarantee the sensing performance in terms of the angle estimation accuracy.
@@
  • The sum rates shown in Fig. 4(a) are obtained based on both the closed-form expressions in Theorem 1 and Monte-Carlo simulations. The analytical and simulation results match well for all considered scenarios, validating our theoretical analysis in Theorem 1.
@@ VI Conclusion
We have investigated mono-static multiuser massive MIMO ISAC systems with linear MRT and ZF precoding. To characterize the system performance and operational properties, we derived closed-form expressions for the achievable communications rate and sensing CRLBs. The analytical findings reveal important properties about ISAC operations in massive MIMO scenarios, such as the mutual interference among the subsystems and the impact of the large number of antennas on communications and sensing performance. We proposed an algorithm for power allocation among the precoders for the communications users and the sensed target to maximize the users’ sum rate with a constraint on the CRLB and power budget. Our theoretical findings and proposed algorithm were verified by extensive numerical results, which show superior communications and sensing performance in massive MIMO ISAC systems. Our future work will consider more challenging sensing use cases such as scenarios with multiple targets located in the near-field region of the massive transmit array.
@@ Appendix A Proof of Lemma 1
$P_{\mathtt{bf}}^{\mathtt{MRT}}=\mathbb{E}\left\{\text{trace}\left({\mathbf{F}}_{\mathtt{MRT}}{\mathbf{F}}_{\mathtt{MRT}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}=\text{trace}\left(\mathbb{E}\left\{{\mathbf{F}}_{\mathtt{MRT}}{\mathbf{F}}_{\mathtt{MRT}}^{\scriptscriptstyle\mathsf{H}}\right\}\right)$
$=\underbrace{\mathbb{E}\left\{\text{trace}\left(\hat{{\mathbf{H}}}\bm{\Gamma}\bm{\Gamma}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}}_{\triangleq\varepsilon_{\mathtt{MRT}}^{\mathtt{tot}}}+\underbrace{\mathbb{E}\left\{\text{trace}\left({\mathbf{a}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\bar{\bm{\eta}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}}_{\triangleq\varepsilon_{0}^{\mathtt{tot}}}.$ (55)
@@
$\varepsilon_{\mathtt{MRT}}^{\mathtt{tot}}$ $=\mathbb{E}\left\{\text{trace}\left(\bm{\Gamma}^{2}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}\right)\right\}=N_{\mathtt{t}}\bm{\xi}_{\mathtt{MRT}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma},$ (56)
where the last equality follows from the fact that the mean of the $k$-th diagonal element of $\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}$ is equal to $N_{\mathtt{t}}\xi_{k}$. Furthermore, noting that ${\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}=N_{\mathtt{t}}$, we have
$\varepsilon_{0}^{\mathtt{tot}}=\mathbb{E}\left\{\text{trace}\left(\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\bar{\bm{\eta}}\right)\right\}=N_{\mathtt{t}}\sum\nolimits_{k=1}^{K}\eta_{k}=N_{\mathtt{t}}\rho.$ (57)
@@
$P_{\mathtt{bf}}^{\mathtt{ZF}}=\mathbb{E}\left\{\text{trace}\left({\mathbf{F}}_{\mathtt{ZF}}{\mathbf{F}}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}=\text{trace}\left(\mathbb{E}\left\{{\mathbf{F}}_{\mathtt{ZF}}{\mathbf{F}}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{H}}\right\}\right)$
$=\text{trace}\left(\mathbb{E}\left\{\left(\hat{{\mathbf{H}}}^{\dagger}\bm{\Gamma}+{\mathbf{a}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\right)\left(\hat{{\mathbf{H}}}^{\dagger}\bm{\Gamma}+{\mathbf{a}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\right)^{\scriptscriptstyle\mathsf{H}}\right\}\right)$
$=\underbrace{\mathbb{E}\left\{\text{trace}\left(\bm{\Gamma}\bm{\Gamma}^{\scriptscriptstyle\mathsf{H}}(\hat{{\mathbf{H}}}^{\dagger})^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}^{\dagger}\right)\right\}}_{\triangleq\varepsilon_{\mathtt{ZF}}^{\mathtt{tot}}}+\underbrace{\mathbb{E}\left\{\text{trace}\left({\mathbf{a}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\bar{\bm{\eta}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}}_{=\varepsilon_{0}^{\mathtt{tot}}}.$ (58)
@@
$\varepsilon_{\mathtt{ZF}}^{\mathtt{tot}}=\mathbb{E}\left\{\text{trace}\left(\bm{\Gamma}^{2}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}\right)\right\}$
$=\!\mathbb{E}\left\{\text{trace}\left(\bm{\Gamma}^{2}\bm{\Xi}^{-1}({\mathbf{Z}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{Z}})^{-1}\right)\right\}\!=\!\sum_{k=1}^{K}\frac{\gamma_{k}}{\xi_{k}}\mathbb{E}\left\{\left[(\hat{{\mathbf{Z}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{Z}}})^{-1}\right]_{kk}\right\}$
$=\!\sum_{k=1}^{K}\!\!\frac{\gamma_{k}}{K\xi_{k}}\mathbb{E}\left\{\text{trace}\left((\hat{{\mathbf{Z}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{Z}}})^{-1}\right)\right\}\!=\!\sum_{k=1}^{K}\frac{\gamma_{k}}{(N_{\mathtt{t}}-K)\xi_{k}},$ (59)
where we have used the property $\mathbb{E}\{\text{trace}((\hat{{\mathbf{Z}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{Z}}})^{-1})\}=\frac{K}{N_{\mathtt{t}}-K}$, since $\hat{{\mathbf{Z}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{Z}}}$ is a central complex Wishart matrix [68]. From (57)–(59), the total power constraint with the ZF precoder is given as (16) where $\bm{\xi}_{\mathtt{bf}}=\bm{\xi}_{\mathtt{ZF}}$ as defined in (17).
@@ Appendix B Proof of Theorem 1
${\mathbf{h}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{f}_{\mathtt{bf}}}_{i}=(\hat{{\mathbf{h}}}_{k}+{\mathbf{e}}_{k})^{\scriptscriptstyle\mathsf{H}}\left(\sqrt{\gamma_{i}}{{\mathbf{w}}_{\mathtt{bf}}}_{i}+\sqrt{\eta_{i}}{\mathbf{a}}\right)$
$=\sqrt{\gamma_{i}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{w}}_{\mathtt{bf}}}_{i}+\sqrt{\eta_{i}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{i}}{{\mathbf{w}}_{\mathtt{bf}}}_{i}+\sqrt{\eta_{i}}{\mathbf{a}}),$ (60)
@@
$\mathbb{E}\left\{\sqrt{\eta_{i}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{i}}{{\mathbf{w}}_{\mathtt{bf}}}_{i}+\sqrt{\eta_{i}}{\mathbf{a}})\right\}=0,\forall k,i$ (61)
as ${\mathbf{e}}_{k}$ and ${{\mathbf{w}}_{\mathtt{bf}}}_{i}$ are independent and they both have zero means. To derive the closed-form SE expressions, in the following we compute $\left|\mathtt{DS}_{k}\right|^{2}$, $\mathbb{E}\{\left|\mathtt{BU}_{k}\right|^{2}\}$, and $\mathbb{E}\{\left|\mathtt{UI}_{kj}\right|^{2}\}$.
@@ B-A MRT Beamforming
$E_{0}$ $=\mathbb{E}\left\{\left|\sqrt{\gamma_{k}}\lVert\hat{{\mathbf{h}}}_{k}\rVert^{2}+\sqrt{\eta_{k}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{k}}\hat{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{\mathbf{a}})\right|^{2}\right\}$
$=\gamma_{k}\mathbb{E}\left\{\lVert\hat{{\mathbf{h}}}_{k}\rVert^{4}\right\}+\eta_{k}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}|^{2}\right\}$
$\hskip 56.9055pt+\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{k}}\hat{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{\mathbf{a}})|^{2}\right\}.$ (64)
Note that the elements of ${\mathbf{a}}$ are deterministic with unit modulus, and $\hat{{\mathbf{h}}}_{k}\sim\mathcal{CN}(0,\xi_{k}\textbf{{I}}_{N_{\mathtt{t}}})$. Thus,
@@
$\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}|^{2}\right\}=\mathbb{E}\left\{\lVert\hat{{\mathbf{h}}}_{k}\rVert^{2}\right\}=N_{\mathtt{t}}\xi_{k},$ (66)
$\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{k}}\hat{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{\mathbf{a}})|^{2}\right\}=N_{\mathtt{t}}\epsilon_{k}\left(\xi_{k}\gamma_{k}+\eta_{k}\right).$ (67)
@@
$=\mathbb{E}\left\{\left|\sqrt{\gamma_{j}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{h}}}_{j}+\sqrt{\eta_{j}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{j}}\hat{{\mathbf{h}}}_{j}+\sqrt{\eta_{j}}{\mathbf{a}})\right|^{2}\right\}$

### v3

@@ III-B Sensing Performance
See Appendix E. $\Box$
We have shown in Remarks 2 and 4 that with the simplified equal power allocation in (23), the interference is mitigated and a very low CRLB can be achieved by a very large $N_{\mathtt{t}}$. However, this does not guarantee a good communications–sensing performance tradeoff. In particular, Remark 1 implies that the massive MIMO ISAC system requires a proper power allocation between communications and sensing; otherwise, there will be significant performance loss due to sensing. In the next section, we present the power allocation problem and its solution.
@@ IV-A Problem Formulation
Note that the proposed optimization in (54) is performed based on the assumption of Remark 3 that the CRLB constraints in (54d) and (54d) are computed without any pointing error. Unlike the sensing performance, there should be little impact on the communication sum-rate objective of (54a) when a small pointing error exists, since it is the presence of the CRLB constraint and not the precise angles associated with the constraint that has the primary effect.
@@ V-C Communications and Sensing Performance
In this section, we demonstrate the communications and sensing performance of MRT and ZF precoding with the proposed power allocation. We first verify the theoretical results for the CRLB in Theorems 1 and Remark 3. In Fig. 4, we compare the CRLB for $\theta$ and $\phi$ obtained by Monte-Carlo simulations based on (39), (40), and the analytical results in Remark 3. We also show the MSE of the MLE in (27) for $\theta$ and $\phi$, which performs an exhaustive search over a fine grid. The sensing SNR is defined as $\frac{P_{\mathtt{t}}L\left|\alpha\right|^{2}}{\sigma_{\mathtt{s}}^{2}}$ [22]. It is observed for both MRT and ZF that the CRLB is a tight lower bound for the MSE of MLE. Therefore, the power allocation constraint on the CRLB threshold can guarantee the sensing performance in terms of the angle estimation accuracy.
@@
  • The sum rates shown in Fig. 5(a) are obtained based on both the closed form expressions in Theorem 1 and Monte-Carlo simulations. The analytical and simulation results match well for all considered scenarios, validating our theoretical analysis in Theorem 1.
@@ VI Conclusion
We have investigated mono-static multiuser massive MIMO ISAC systems with linear MRT and ZF precoding. To characterize the system performance and operational properties, we derived closed form expressions for the achievable communications rate and sensing CRLBs. The analytical findings reveal important properties about ISAC operations in massive MIMO scenarios, such as the mutual interference among the subsystems and the impact of the large number of antennas on communications and sensing performance. We proposed an algorithm for power allocation among the precoders for the communications users and the sensed target to maximize the users’ sum rate with a constraint on the CRLB and power budget. Our theoretical findings and proposed algorithm were verified by extensive numerical results, which show superior communications and sensing performance in massive MIMO ISAC systems.
@@ Appendix A Proof of Lemma 1
$P_{\mathtt{bf}}^{\mathtt{MRT}}=\mathbb{E}\left\{\mathtt{trace}\left({\mathbf{F}}_{\mathtt{MRT}}{\mathbf{F}}_{\mathtt{MRT}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}=\mathtt{trace}\left(\mathbb{E}\left\{{\mathbf{F}}_{\mathtt{MRT}}{\mathbf{F}}_{\mathtt{MRT}}^{\scriptscriptstyle\mathsf{H}}\right\}\right)$
$=\text{trace}\left(\mathbb{E}\left\{\left(\hat{{\mathbf{H}}}\bm{\Gamma}+{{\mathbf{v}}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\right)\left(\hat{{\mathbf{H}}}\bm{\Gamma}+{{\mathbf{v}}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\right)^{\scriptscriptstyle\mathsf{H}}\right\}\right)$
$=\underbrace{\mathbb{E}\left\{\mathtt{trace}\left(\hat{{\mathbf{H}}}\bm{\Gamma}\bm{\Gamma}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}}_{\triangleq\varepsilon_{\mathtt{MRT}}^{\mathtt{tot}}}+\underbrace{\mathbb{E}\left\{\mathtt{trace}\left({{\mathbf{v}}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\bar{\bm{\eta}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}}_{\triangleq\varepsilon_{0}^{\mathtt{tot}}}.$ (62)
@@
$\varepsilon_{\mathtt{MRT}}^{\mathtt{tot}}$ $=\mathbb{E}\left\{\mathtt{trace}\left(\bm{\Gamma}^{2}\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}\right)\right\}=N_{\mathtt{t}}\bm{\xi}_{\mathtt{MRT}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma},$ (63)
where the last equality follows from the fact that the mean of the $k$-th diagonal element of $\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}$ is equal to $N_{\mathtt{t}}\xi_{k}$. Furthermore, recall that ${{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}=N_{\mathtt{t}}$, we have
$\varepsilon_{0}^{\mathtt{tot}}=\mathbb{E}\left\{\mathtt{trace}\left(\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\bar{\bm{\eta}}\right)\right\}=N_{\mathtt{t}}\sum\nolimits_{k=1}^{K}\eta_{k}=N_{\mathtt{t}}\rho.$ (64)
@@
$P_{\mathtt{bf}}^{\mathtt{ZF}}=\mathbb{E}\left\{\mathtt{trace}\left({\mathbf{F}}_{\mathtt{ZF}}{\mathbf{F}}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}=\mathtt{trace}\left(\mathbb{E}\left\{{\mathbf{F}}_{\mathtt{ZF}}{\mathbf{F}}_{\mathtt{ZF}}^{\scriptscriptstyle\mathsf{H}}\right\}\right)$
$=\text{trace}\left(\mathbb{E}\left\{\left(\hat{{\mathbf{H}}}^{\dagger}\bm{\Gamma}+{{\mathbf{v}}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\right)\left(\hat{{\mathbf{H}}}^{\dagger}\bm{\Gamma}+{{\mathbf{v}}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\right)^{\scriptscriptstyle\mathsf{H}}\right\}\right)$
$=\underbrace{\mathbb{E}\left\{\mathtt{trace}\left(\bm{\Gamma}\bm{\Gamma}^{\scriptscriptstyle\mathsf{H}}(\hat{{\mathbf{H}}}^{\dagger})^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}}^{\dagger}\right)\right\}}_{\triangleq\varepsilon_{\mathtt{ZF}}^{\mathtt{tot}}}+\underbrace{\mathbb{E}\left\{\mathtt{trace}\left({{\mathbf{v}}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}}\bar{\bm{\eta}}{{\mathbf{v}}}^{\scriptscriptstyle\mathsf{H}}\right)\right\}}_{=\varepsilon_{0}^{\mathtt{tot}}}.$ (65)
@@
$\varepsilon_{\mathtt{ZF}}^{\mathtt{tot}}=\mathbb{E}\left\{\mathtt{trace}\left(\bm{\Gamma}^{2}(\hat{{\mathbf{H}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{H}}})^{-1}\right)\right\}$
$=\!\mathbb{E}\left\{\mathtt{trace}\left(\bm{\Gamma}^{2}\bm{\Xi}^{-1}({\mathbf{Z}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{Z}})^{-1}\right)\right\}\!=\!\sum_{k=1}^{K}\frac{\gamma_{k}}{\xi_{k}}\mathbb{E}\left\{\left[(\hat{{\mathbf{Z}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{Z}}})^{-1}\right]_{kk}\right\}$
$=\!\sum_{k=1}^{K}\!\!\frac{\gamma_{k}}{K\xi_{k}}\mathbb{E}\left\{\mathtt{trace}\left((\hat{{\mathbf{Z}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{Z}}})^{-1}\right)\right\}\!=\!\sum_{k=1}^{K}\frac{\gamma_{k}}{(N_{\mathtt{t}}-K)\xi_{k}},$ (66)
where we have used the property $\mathbb{E}\{\mathtt{trace}((\hat{{\mathbf{Z}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{Z}}})^{-1})\}=\frac{K}{N_{\mathtt{t}}-K}$, since $\hat{{\mathbf{Z}}}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{Z}}}$ is a central complex Wishart matrix [68]. From (64)–(66), the total power constraint with the ZF precoder is given as (16) where $\bm{\xi}_{\mathtt{bf}}=\bm{\xi}_{\mathtt{ZF}}$ as defined in (17).
@@ Appendix B Proof of Theorem 1
${\mathbf{h}}_{k}^{\scriptscriptstyle\mathsf{H}}{\mathbf{f}_{\mathtt{bf}}}_{i}=(\hat{{\mathbf{h}}}_{k}+{\mathbf{e}}_{k})^{\scriptscriptstyle\mathsf{H}}\left(\sqrt{\gamma_{i}}{{\mathbf{w}}_{\mathtt{bf}}}_{i}+\sqrt{\eta_{i}}{{\mathbf{v}}}\right)$
$=\sqrt{\gamma_{i}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{w}}_{\mathtt{bf}}}_{i}+\sqrt{\eta_{i}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{i}}{{\mathbf{w}}_{\mathtt{bf}}}_{i}+\sqrt{\eta_{i}}{{\mathbf{v}}}),$ (67)
@@
$\mathbb{E}\left\{\sqrt{\eta_{i}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{i}}{{\mathbf{w}}_{\mathtt{bf}}}_{i}+\sqrt{\eta_{i}}{{\mathbf{v}}})\right\}=0,\forall k,i$ (68)
as ${\mathbf{e}}_{k}$ and ${{\mathbf{w}}_{\mathtt{bf}}}_{i}$ are independent and they both have zero means. To derive the closed form SE expressions, in the following we compute $\left|\mathtt{DS}_{k}\right|^{2}$, $\mathbb{E}\{\left|\mathtt{BU}_{k}\right|^{2}\}$, and $\mathbb{E}\{\left|\mathtt{UI}_{kj}\right|^{2}\}$.
@@ B-A MRT Beamforming
$E_{0}$ $=\mathbb{E}\left\{\left|\sqrt{\gamma_{k}}\lVert\hat{{\mathbf{h}}}_{k}\rVert^{2}+\sqrt{\eta_{k}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{k}}\hat{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{{\mathbf{v}}})\right|^{2}\right\}$
$=\gamma_{k}\mathbb{E}\left\{\lVert\hat{{\mathbf{h}}}_{k}\rVert^{4}\right\}+\eta_{k}\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}|^{2}\right\}$
$\hskip 56.9055pt+\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{k}}\hat{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{{\mathbf{v}}})|^{2}\right\}.$ (71)
Note that with ${\mathbf{v}}={\mathbf{a}}(\tilde{\theta},\tilde{\phi})$, the elements of ${{\mathbf{v}}}$ are deterministic with unit modulus, and $\hat{{\mathbf{h}}}_{k}\sim\mathcal{CN}(0,\xi_{k}\textbf{{I}}_{N_{\mathtt{t}}})$. Thus,
@@
$\mathbb{E}\left\{|\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}|^{2}\right\}=\mathbb{E}\left\{\lVert\hat{{\mathbf{h}}}_{k}\rVert^{2}\right\}=N_{\mathtt{t}}\xi_{k},$ (73)
$\mathbb{E}\left\{|{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{k}}\hat{{\mathbf{h}}}_{k}+\sqrt{\eta_{k}}{{\mathbf{v}}})|^{2}\right\}=N_{\mathtt{t}}\epsilon_{k}\left(\xi_{k}\gamma_{k}+\eta_{k}\right).$ (74)
@@
$=\mathbb{E}\left\{\left|\sqrt{\gamma_{j}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}\hat{{\mathbf{h}}}_{j}+\sqrt{\eta_{j}}\hat{{\mathbf{h}}}_{k}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}}+{\mathbf{e}}_{k}^{\scriptscriptstyle\mathsf{H}}(\sqrt{\gamma_{j}}\hat{{\mathbf{h}}}_{j}+\sqrt{\eta_{j}}{{\mathbf{v}}})\right|^{2}\right\}$